### v3

@@ 2 Problem Formulation
[…]
@@ 3.3 Filtering, Optimization and Results Printing (lines 58-81 and and 91-122 )
With rfill, a loop is run to identify neighboring elements within the filter radius, and their indices are recorded in I (line 68). Concurrently, a corresponding index array J is created (line 69). The $x,\,y,$ and distance information of neighboring elements, along with their respective indices, are consolidated into arrays newx, newy, and newz, respectively. Arrays are combined to form the matrix DD (lines 71-74) containing the $x$ and $y$ coordinates and distance of neighboring FEs. The filtering matrix HHs is evaluated on lines 76-77 using spdiags and coo_matrix SciPy that create a sparse diagonal matrix with indices based on the distances in DD.

### v4

@@ Abstract
Published [^1] This pdf is the personal version of an article whose final publication is available at Advances in Multidisciplinary Design, Analysis and Optimization   in Advances in Multidisciplinary Design, Analysis and Optimization, DOI:10.1007/978-981-96-1158-4_35
@@ 2 Problem Formulation
The design domain is parameterized using Nelem hexagonal finite elements (FEs). Each element is attached with a design variable $\rho_{j}\in[0,\,1]$. We use the modified SIMP [2], wherein the stiffness matrix of element $j$ is determined as
@@ 3.3 Filtering, Optimization and Results Printing (lines 58-81 and and 91-122 )
With rfill, a loop is run to identify neighboring elements within the filter radius, and their indices are recorded in I (line 68). Concurrently, a corresponding index array J is created (line 69). The x,y,x,\,y, and distance information of neighboring elements, along with their respective indices, are consolidated into arrays newx, newy, and newz, respectively. Arrays are combined to form the matrix DD (lines 71-74) containing the $x$ and $y$ coordinates and distance of neighboring FEs. The filtering matrix HHs is evaluated on lines 76-77 using spdiags and coo_matrix SciPy that create a sparse diagonal matrix with indices based on the distances in DD.